RC meeting and Run Plan, Tuesday September 27th

From SBS wiki
Jump to: navigation, search


RC Meeting Information

Run Coordinator: Donald Jones

Location: Counting House Conference Room

Time: 3:00PM

Commissioning run plans

  • Beam line (1-2 shifts) Sept 30.
    • Beam centering and target alignment
    • Raster calibration
    • BCM and BPM calibration
  • Corrector adjustment procedure
  • GEMs run plan
  • Charge feedback system
  • Optics and calibration
  • Moller polarimeter (1 shift) Oct 7
  • Target commissioning
  • Bigbite commissioning (preshower, shower, GRINCH)
  • HCal commissioning

Subsystem status reports

  • Hall A Status -- Jessie Butler
  • Target status/timeline (training?) -- Arun
  • Beam line:
    • Monitors (calibration) -- Simona
    • Correctors -- Bogdan, Iuliia
  • Tentative run plan document -- Bogdan
  • Alarms
  • Analyzer -- Andrew
  • DAQ -- Alex
  • HCal -- Jiwan
  • Spectrometer magnets -- Jack
  • GEMs/Tracking -- Holly
    • Are new HV controls, alarms, gas flow monitoring procedures, EPCIS tracking in place?
    • Shift checklist ok?
  • Shower Calorimeter -- Provakar
  • Beam polarization -- Don


  • RC reported that magnetic field measurements were ongoing to see if fields and gradients are sufficiently low inside the target area. Steel plates were installed partially over SBS window.
  • Mark reported that Ed Folts and Patricia require that the experiment be ready including the full target before signing off on a run permit. That means we need the target installed and surveyed, the laser aligned, properly focused, polarization measured and tested at full power.
  • Arun reported that measurements had been done again at full BB and SBS magnet currents and although the Hall probes read values like 40 and 50, he was able to change the values by something like 40 just by turning on the large Helmholtz coils. These coils max out at something like 25 G, so we know that the probe readings are not calibrated in Gauss. We now have data in hand which need to be analyzed and a rough calibration placed on the probe readings to get a better idea of what the fields actually are.
  • Gordon reported real progress on laser alignment and characterization. Work will continue tonight.
  • After discussions between Arun, Jessie, Todd and Gordon we concluded that the final plan would be to take the Hall into CANS mode this evening and work on the laser until 10PM. At 10PM, Arun and Todd will work to install the target ladder, including optics foils, Ukraine and the reference cell. Todd will do the final gas line installation. The oven will then be heated up in time for survey in the morning. At 7:30AM the survey and alignment folks will survey the target ladder position and verify the motion is sufficiently repeatable. While this is ongoing, Mark will work to get the OSP for field measurements approved with the hope that it is completely signed off by 2-3PM when the field measurements will resume under Jack and Arun. We are aiming still to get beam Friday if possible.